# Test-Time Compute: from System-1 Thinking to System-2 Thinking

### v2

@@ 4.2.1 Repeated Sampling
Verification strategies include two types: majority voting and best-of-N (BoN) sampling. Majority voting (Li et al., 2024c; Lin et al., 2024a) selects the most frequently occurring answer in the samples as the final answer, which is motivated by ensemble learning. Majority voting is simple yet effective. For instance, self-consistency CoT (Wang et al., 2023d) can improve accuracy by 18% over vanilla CoT in math reasoning tasks. However, the majority does not always hold the truth, as they may make similar mistakes. Therefore, some studies perform validation and filtering before voting. For example, the PROVE framework (Toh et al., 2024) converts CoT into executable programs, filtering out samples if the program’s results are inconsistent with the reasoning chain’s outcomes.
Best-of-N sampling uses a verifier to score each response and selects the one with the highest score as the final answer (Stiennon et al., 2020; Cobbe et al., 2021; Nakano et al., 2022). Li et al. (2023c) propose a voting-based BoN variant, which performs weighted voting on all answers based on the verifier’s scores and selects the answer with the highest score. (Liu et al., 2025a) design BoN in a knockout tournament, using pairwise comparison verifiers to filter out the best response. In addition, some works aim to improve the efficiency of BoN. Inspired by speculative decoding, Zhang et al. (2024i); Qiu et al. (2024a); Sun et al. (2024) and Manvi et al. (2024) evaluate each reasoning step and prune low-scoring sampled results, halting further generation for those paths, thereby significantly reducing the overall time cost. PRS (Ye and Ng, 2024) enables LLMs to self-critique and self-correct, guiding the model to generate expected responses with fewer sampling times.
@@
Repeated sampling, especially the BoN strategy, has proven to be a simple yet effective method, even surpassing models fine-tuned with RLHF (Gao et al., 2023a; Hou et al., 2024). However, it comes at the cost of inference times that are difficult to afford in practical applications. Therefore, many studies have attempted to train the model by BoN sampling to approximate the BoN distribution, thereby reducing the search space during inference. ReST (Gulcehre et al., 2023) samples responses with reward values above a threshold from the policy model as self-training data and fine-tune the policy model by offline reinforcement learning. In each iteration, ReST samples new training data. vBoN (Amini et al., 2024), BoNBoN (Gui et al., 2024) and BOND (Sessa et al., 2024) derive the BoN distribution and minimize the difference between the policy model’s distribution and the BoN distribution. Chow et al. (2024) design a BoN-aware loss to make the policy model more exploratory during fine-tuning.
4.2.2 Self-correction
@@
The feedback used for self-correction is typically presented in natural language and comes from various sources, including human evaluation, tool checking, external model evaluation, and intrinsic feedback. Human evaluation is the gold standard for feedback, but due to its high cost and limited scalability, it is mainly used in early research to explore the upper limits of self-correction capabilities (Tandon et al., 2021; Elgohary et al., 2021; Tandon et al., 2022). For certain domain-specific tasks, external tool checking provides accurate and efficient feedback (Gou et al., 2024; Chen et al., 2024d; Gao et al., 2023b). For example, Yasunaga and Liang (2020) propose to obtain feedback from compilers in code repair and generation tasks. In embodied tasks, the environment can provide precise feedback on the action trajectories of LLM-based agents (Wang et al., 2024b).
@@
External model evaluation is an effective feedback source for general tasks, such as various verbal-based critique models described in Section 4.1. For example, Paul et al. (2024) first define multiple error types for natural language reasoning tasks and then design the corresponding feedback templates. They train an evaluation model using synthetic feedback training data, and with the critic, the reasoning model achieves substantial performance improvement. Multi-agent debate (Du et al., 2023b; Xiong et al., 2023; Liang et al., 2024b; Chen et al., 2024b; Wang et al., 2024g) is another mechanism that leverages external feedback to enhance reasoning capabilities. In this approach, models do not have distinct roles as reasoners and critics. Instead, multiple models independently conduct reasoning, critique each other, and defend or refine their reasoning based on feedback. This process continues until agents reach a consensus or a judge model summarizes the final reasoning results. The multi-agent debate has shown its potential in fact-checking (Kim et al., 2024a; Khan et al., 2024), commonsense QA (Xiong et al., 2023), faithful evaluations (Chan et al., 2024), and complex reasoning (Du et al., 2023b; Cheng et al., 2024). However, multi-agent debate may be unstable, as LLMs are susceptible to adversarial information and may revise correct answers to incorrect ones in response to misleading inputs (Laban et al., 2024; Amayuelas et al., 2024). Therefore, a successful multi-agent debate requires that LLMs maintain their stance when faced with incorrect answers from other models while remaining open to valid suggestions (Stengel-Eskin et al., 2024). In general, the more LLMs involved in the debate, the stronger the overall reasoning performance. However, this significantly increases the number of LLM inferences required, and the length of input context, posing a major challenge to LLM inference costs (Liu et al., 2024c). To reduce debate inference costs, Li et al. (2024g) investigate the impact of topological connections among multiple agents and show that sparse connections, such as ring structures, are not inferior to the fully connected topology. GroupDebate (Liu et al., 2024c) divides LLMs into groups that conduct debates internally and only share the consensus results between groups.
@@
Self-critique assumes that LLMs can self-evaluate their outputs and optimize them through intrinsic feedback (Yuan et al., 2024c). This idea stems from a fundamental principle in computational complexity theory: verifying whether a solution is correct is typically easier than solving the problem. Bai et al. (2022) propose self-correcting harmful responses from LLMs by prompting themselves. Self-Refine (Madaan et al., 2023) and RCI Prompting (Kim et al., 2023) iteratively prompt LLMs to self-correct their responses in tasks such as arithmetic reasoning. IoE (Li et al., 2024e) observes that LLMs may over-criticize themselves during self-critique, leading to performance degradation, and designs prompt to guide LLMs in assessing confidence. ProgCo (Song et al., 2025b) leverage the advantages of code in expressing complex logic, enabling LLMs to generate responses in pseudo-code form, followed by self-critique and refinement. SETS (Chen et al., 2025a) combines the strengths of repeated sampling and self-critique, applying self-critique and correction to each sampled reasoning path and choosing the final solution via majority voting.
@@
The effectiveness of self-correction, especially the self-critique, has remained controversial. Several empirical studies on code generation (Olausson et al., 2024), commonsense QA (Huang et al., 2024a), math problem-solving (Wang et al., 2024f), planning (Valmeekam et al., 2023a), and graph coloring (Stechly et al., 2023) confirm that self-correction is not a guaranteed solution for improving performance. Kamoi et al. (2024) think the effectiveness of self-correction has been overestimated. Previous successes either rely on oracle answers or weak initial answers. Only tasks that can be broken down into easily verifiable sub-tasks can truly benefit from self-correction. They suggest fine-tuning specific evaluation models to achieve better self-correction. Zhang et al. (2024h) try to interpret and alleviate the failure of self-critique via human-like cognitive bias. Tyen et al. (2024) decouple the abilities of LLMs to identify and correct errors and create the corresponding evaluation datasets. The evaluation results show that LLMs do not lack the ability to correct errors during self-correction, and their main performance bottleneck lies in locating the errors. Yang et al. (2024b) decompose self-critique into confidence and critique capabilities. Empirical studies show that fine-tuning is necessary to enhance both capabilities simultaneously, while prompt engineering can only achieve a trade-off.
@@
Most of the aforementioned self-correction methods demonstrate significant performance improvements on advanced closed-source large models or open-source LLMs with over 70B parameters. However, for medium-scale open-source models with weaker capabilities, we need to further fine-tune them to unlock their self-correction capabilities. Supervised fine-tuning optimizes the model using high-quality multi-turn correction data, either manually annotated (Saunders et al., 2022), self-rationalize (Zelikman et al., 2022; Yuan et al., 2025b), multi-agent debate (Subramaniam et al., 2025) or sampled from stronger LLMs (An et al., 2023; Paul et al., 2024; Qu et al., 2024; Gao et al., 2024c; Zhang et al., 2024k; Xi et al., 2024). GLoRe (Havrilla et al., 2024) considers that LLMs need global or local refinement for different types of errors. To address this, they construct training sets for global and local refinement, train verifiers to identify global and local errors, and develop LLMs for refinement based on different global or local feedback signals. Xi et al. (2024) design a scalable framework for synthesizing self-correction training data, enabling reasoning models to generate controlled errors and receive feedback from critics to self-correct. Although SFT is effective, training data from offline-generated self-correction trajectories can only simulate limited correction patterns. This leads to the distribution mismatch with the actual self-correction behavior during model inference. Self-correct (Welleck et al., 2023) adopts online imitation learning, re-sampling new self-correction trajectories for training after each training epoch. To further expand the exploration space of LLMs, many studies adopt flexible RL algorithms to surpass the performance limits of SFT. SCoRe (Kumar et al., 2024) proposes using the multi-turn RL method to improve self-critique and self-correction capability. T1 (Hou et al., 2025) employs self-correction training data for SFT cold-start, followed by RL training using the RLOO algorithm (Ahmadian et al., 2024). During the RL phase, high-temperature sampling and entropy rewards encourage the LLM to explore more diverse reasoning paths. Deepseek-R1 (Guo et al., 2025a) uses rule-based rewards and the GRPO algorithm (Shao et al., 2024) for RL training. It also demonstrates RL’s immense potential, even without SFT cold-start, its exploration capabilities suffice to endow LLMs with strong reasoning abilities.
4.2.3 Tree Searching
@@
In LLM reasoning, current search algorithms include uninformed search and heuristic search. Uninformed search explores the search space according to a fixed rule. For example, tree-of-thought (ToT) (Yao et al., 2023) adopts the BFS or DFS to search, while Xie et al. (2023) use beam search. Uninformed search is usually less efficient for problems with large search spaces, so heuristic search strategies represented by A (Meng et al., 2024; Wang et al., 2024a) and MCTS (Hao et al., 2023; Bi et al., 2024; Park et al., 2024) are widely used in reasoning tasks. MCTS, which eliminates the need for explicit heuristics, leverages stochastic simulations and adaptive tree expansion under uncertain environments, making it well-suited for large state spaces. It optimizes search results gradually through four steps: selection, expansion, simulation, and backpropagation, approaching the optimal solution. In contrast, A uses a heuristic function-guided deterministic search to guarantee optimal paths, but its performance depends on the design of the heuristic function. As a result, MCTS has been successfully applied to tasks such as RAG (Hu et al., 2024b; Jiang et al., 2024; Li and Ng, 2024; Feng et al., 2025), QA (Luo et al., 2025a; Gan et al., 2025), hallucinations mitigation (Cheng et al., 2025), text-to-SQL (Yuan et al., 2025a), etc. Additionally, Long (2023) trains an LLM controller using reinforcement learning to guide the LLM reasoner’s search path, and Chari et al. (2025) utilizes ant colony evolutionary algorithm to guide tree search.
@@
The value function evaluates the value of each state and guides the tree to expand towards branches with higher values in heuristic tree search. Xu (2023) train an energy function by noise-contrastive estimation as the value function. RAP (Hao et al., 2023) designs a series of heuristic value functions, including the likelihood of the action, the confidence of the state, self-evaluation results, and task-specific reward, and combines them according to task requirements. Reliable and generalized value functions facilitate the application of MCTS to more complex problems with deeper search spaces. AlphaMath (Chen et al., 2024a) and TS-LLM (Feng et al., 2024b) replace the hand-crafted value function with a learned LLM value function, automatically generating reasoning process and step-level evaluation signals in MCTS. VerifierQ (Qi et al., 2024a) integrates implicit Q-learning and contrastive Q-learning to train the value function, effectively mitigating the overestimation issue at the step level. Traditional MCTS methods expand only one trajectory, while rStar (Qi et al., 2024b) argues that the current value function struggles to guide the selection of the optimal path accurately. Therefore, rStar retains multiple candidate paths and performs reasoning with another LLM, ultimately selecting the path where both LLMs’ reasoning results are consistent. Gao et al. (2024d) propose SC-MCTS which combines multiple reward models, including contrastive reward, likelihood and self-evaluation as value functions. MCTSr (Zhang et al., 2024b) and SR-MCTS (Zhang et al., 2024c) take complete responses as nodes, expanding the search space through self-critique and correction. SR-MCTS utilizes pairwise preference rewards and global quantile score as the value function, offering more robust value function estimation compared to step-based MCTS.
@@
Tree search can guide LLMs to generate long reasoning processes, and these data help train LLMs with stronger reasoning abilities (Zhai et al., 2024; Xu et al., 2024a; Guan et al., 2025). ReST-MCTS* (Zhang et al., 2024a) uses process rewards as a value function to guide MCTS, collecting high-quality reasoning trajectories and the value of each step to improve the policy model and reward model. Due to the step-by-step exploration of tree search, it can obtain finer-grained step-level feedback signals. MCTS-DPO (Xie et al., 2024) collects step-level preference data through MCTS and uses DPO for preference learning. AlphaLLM-CPL (Wang et al., 2024k) ranks trajectories based on preference reward gaps and policy prediction gaps, employing curriculum learning to efficiently utilize MCTS-collected trajectories. Recently, many o1-like technical reports (Qin et al., 2024b; Zhao et al., 2024b; Zhang et al., 2024j) have also confirmed the necessity of using tree search to construct high-quality long reasoning chain data for training.
@@
5 Future Directions
5.1 Generalizable System-2 Model

### v3

@@ 4.2.1. Repeated Sampling
Verification strategies include two types: majority voting and best-of-N (BoN) sampling. Majority voting (Li et al., 2024h; Lin et al., 2024a) selects the most frequently occurring answer in the samples as the final answer, which is motivated by ensemble learning. Majority voting is simple yet effective. For instance, self-consistency CoT (Wang et al., 2023b) can improve accuracy by 18% over vanilla CoT in math reasoning tasks. However, the majority does not always hold the truth, as they may make similar mistakes. Therefore, some studies perform filtering or weighting before voting. For example, the PROVE framework (Toh et al., 2024) converts CoT into executable programs, filtering out samples if the program’s results are inconsistent with the reasoning chain’s outcomes. Kang et al. (2025b) propose the self-certainty metric to weight votes based on their ranking. Huang et al. (2025a) use the calibrated confidence as vote weights, and adjust the number of samples based on response agreement to improve efficiency (Aggarwal et al., 2023). RPC (Zhou et al., 2025) demonstrates theoretically and experimentally that filtering low-probability samples improves the performance and efficiency of majority voting.
Best-of-N sampling uses a verifier to score each response and selects the one with the highest score as the final answer (Stiennon et al., 2020; Cobbe et al., 2021; Nakano et al., 2022; Liu et al., 2025a). Li et al. (2023a) propose a voting-based BoN variant, which performs weighted voting on all answers based on the verifier’s scores and selects the answer with the highest score. (Liu et al., 2025f) design BoN in a knockout tournament, using pairwise comparison verifiers to filter out the best response. In addition, some works aim to improve the efficiency of BoN. Inspired by speculative decoding, Zhang et al. (2024b); Qiu et al. (2024a); Sun et al. (2024); Yu et al. (2025a) and Manvi et al. (2024) evaluate each reasoning step and prune low-scoring sampled results, halting further generation for those paths, thereby significantly reducing the overall time cost. PRS (Ye and Ng, 2024) enables LLMs to self-critique and self-correct, guiding the model to generate expected responses with fewer sampling times. Li et al. (2025a) compare BoN and majority voting, demonstrating that BoN is suitable for harder questions with moderate diversity of response distribution.
@@
Repeated sampling has proven to be a simple yet effective method, even surpassing models fine-tuned with RLHF (Gao et al., 2023b; Hou et al., 2024). However, it costs much inference time that is difficult to afford in practical applications. Therefore, many studies have attempted to train the model by BoN sampling to approximate the BoN distribution, thereby reducing the search space during inference. ReST (Gulcehre et al., 2023) samples responses with reward values above a threshold from the policy model as self-training data and fine-tunes the policy model by offline reinforcement learning. In each iteration, ReST samples new training data. vBoN (Amini et al., 2024), BoNBoN (Gui et al., 2024) and BOND (Sessa et al., 2024) derive the BoN distribution and minimize the difference between the policy model’s distribution and the BoN distribution. Chow et al. (2024) design a BoN-aware loss to make the policy model more exploratory during fine-tuning.
4.2.2. Self-correction
@@
Figure 5. Illustration of self-correction with tool checking, external model evaluation, and self-critique.
@@
The feedback used for self-correction is typically presented in natural language and comes from various sources, including human evaluation, tool checking, external model evaluation, and intrinsic feedback. Human evaluation is the gold standard for feedback, but due to its high cost and limited scalability, it is mainly used in early research to explore the upper limits of self-correction capabilities (Tandon et al., 2021; Elgohary et al., 2021; Tandon et al., 2022). For certain domain-specific tasks, external tool checking provides accurate and efficient feedback (Gou et al., 2024; Chen et al., 2024c; Gao et al., 2023a). For example, Yasunaga and Liang (2020) propose to obtain feedback from compilers in code repair and generation tasks. In embodied tasks, the environment can provide precise feedback on the action trajectories of LLM-based agents (Wang et al., 2024e).
@@
External model evaluation is an effective feedback source for general tasks, such as various verbal-based critique models described in Section 4.1. For example, Paul et al. (2024) first define multiple error types for natural language reasoning tasks and then design the corresponding feedback templates. They train an evaluation model using synthetic feedback training data, and with the critic, the reasoning model achieves substantial performance improvement. Multi-agent debate (Du et al., 2023b; Xiong et al., 2023; Liang et al., 2024a; Chen et al., 2024e; Wang et al., 2024l) is another mechanism that leverages external feedback to enhance reasoning capabilities. In this approach, models do not have distinct roles as reasoners and critics. Instead, multiple models independently conduct reasoning, critique each other, and defend or refine their reasoning based on feedback. This process continues until agents reach a consensus or a judge model summarizes the final reasoning results. The multi-agent debate has shown its potential in fact-checking (Kim et al., 2024b; Khan et al., 2024), commonsense QA (Xiong et al., 2023), faithful evaluations (Chan et al., 2024), and complex reasoning (Du et al., 2023b). However, multi-agent debate may be unstable, as LLMs are susceptible to adversarial information and may revise correct answers to incorrect ones in response to misleading inputs (Laban et al., 2024; Amayuelas et al., 2024). Therefore, a successful multi-agent debate requires that LLMs maintain their stance when faced with incorrect answers from other models while remaining open to valid suggestions (Stengel-Eskin et al., 2024). In general, the more LLMs involved in the debate, the stronger the overall reasoning performance. However, this significantly increases the number of LLM inferences required, and the length of input context, posing a major challenge to LLM inference costs (Liu et al., 2024b). To reduce debate inference costs, Li et al. (2024b) investigate the impact of topological connections among multiple agents and show that sparse connections, such as ring structures, are not inferior to the fully connected topology. GroupDebate (Liu et al., 2024b) divides LLMs into groups that conduct debates internally and only share the consensus results between groups.
@@
Self-critique assumes that LLMs can self-evaluate their outputs and optimize them through intrinsic feedback (Yuan et al., 2024c). This idea stems from a fundamental principle in computational complexity theory: verifying whether a solution is correct is typically easier than solving the problem. Bai et al. (2022) propose self-correcting harmful responses from LLMs by prompting themselves. Self-Refine (Madaan et al., 2023) and RCI Prompting (Kim et al., 2023) iteratively prompt LLMs to self-correct their responses in tasks such as arithmetic reasoning. IoE (Li et al., 2024a) observes that LLMs may over-criticize themselves during self-critique, leading to performance degradation, and designs prompt to guide LLMs in assessing confidence. ProgCo (Song et al., 2025b) leverage the advantages of code in expressing complex logic, enabling LLMs to generate responses in pseudo-code form, followed by self-critique and refinement. SETS (Chen et al., 2025c) combines the strengths of repeated sampling and self-critique, applying self-critique and correction to each sampled reasoning path and choosing the final solution via majority voting. S1 (Muennighoff et al., 2025) adds the “wait” to the reasoning process, prompting the LLM to critique its reasoning.
@@
The effectiveness of self-correction, especially the self-critique, has remained controversial. Several empirical studies on code generation (Olausson et al., 2024), commonsense QA (Huang et al., 2024a), math problem-solving (Wang et al., 2024c), planning (Valmeekam et al., 2023a), and graph coloring (Stechly et al., 2023) confirm that self-correction is not a guaranteed solution for improving performance. Kamoi et al. (2024) think the effectiveness of self-correction has been overestimated. Previous successes either rely on oracle answers or weak initial answers. Only tasks that can be broken down into easily verifiable sub-tasks can truly benefit from self-correction. They suggest fine-tuning specific evaluation models to achieve better self-correction. Zhang et al. (2024e) try to interpret and alleviate the failure of self-critique via human-like cognitive bias. Tyen et al. (2024) decouple the abilities of LLMs to identify and correct errors and create the corresponding evaluation datasets. The evaluation results show that LLMs do not lack the ability to correct errors during self-correction, and their main performance bottleneck lies in locating the errors. Yang et al. (2024b) decompose self-critique into confidence and critique capabilities. Empirical studies show that fine-tuning is necessary to enhance both capabilities simultaneously, while prompt engineering can only achieve a trade-off.
@@
Most of the aforementioned self-correction methods demonstrate significant performance improvements on advanced closed-source large models or open-source LLMs with over 70B parameters. However, for medium-scale open-source models with weaker capabilities, we need to further fine-tune them to unlock their self-correction capabilities. Supervised fine-tuning optimizes the model using high-quality multi-turn correction data, either manually annotated (Saunders et al., 2022), self-rationalize (Zelikman et al., 2022; Yuan et al., 2025a), multi-agent debate (Subramaniam et al., 2025) or sampled from stronger LLMs (An et al., 2023; Paul et al., 2024; Qu et al., 2024; Gao et al., 2024a; Zhang et al., 2024a; Xi et al., 2024). GLoRe (Havrilla et al., 2024) considers that LLMs need global or local refinement for different types of errors. To address this, they construct training sets for global and local refinement, train verifiers to identify global and local errors, and develop LLMs for refinement based on different global or local feedback signals. Xi et al. (2024) design a scalable framework for synthesizing self-correction training data, enabling reasoning models to generate controlled errors and receive feedback from critics to self-correct. Although SFT is effective, training data from offline-generated self-correction trajectories can only simulate limited correction patterns. This leads to the distribution mismatch with the actual self-correction behavior during model inference. Self-correct (Welleck et al., 2023) adopts online imitation learning, re-sampling new self-correction trajectories for training after each training epoch. To further expand the exploration space of LLMs, many studies adopt flexible RL algorithms to surpass the performance limits of SFT. SCoRe (Kumar et al., 2024) proposes using the multi-turn RL method to improve self-critique and self-correction capability. T1 (Hou et al., 2025) employs self-correction training data for SFT cold-start, followed by RL training using the RLOO algorithm (Ahmadian et al., 2024). During the RL phase, high-temperature sampling and entropy rewards encourage the LLM to explore more diverse reasoning paths. Deepseek-R1 (Guo et al., 2025a) uses rule-based rewards and the GRPO algorithm (Shao et al., 2024) for RL training. It also demonstrates RL’s immense potential, even without SFT cold-start, its exploration capabilities suffice to endow LLMs with strong reasoning abilities.
4.2.3. Tree Searching
@@
In LLM reasoning, current search algorithms include uninformed search and heuristic search. Uninformed search explores the search space according to a fixed rule. For example, tree-of-thought (ToT) (Yao et al., 2023) adopts the BFS or DFS to search, while Xie et al. (2023) use beam search. Uninformed search is usually less efficient for problems with large search spaces, so heuristic search strategies represented by A (Meng et al., 2024; Wang et al., 2024a) and MCTS (Hao et al., 2023; Bi et al., 2024; Park et al., 2024) are widely used in reasoning tasks. MCTS, which eliminates the need for explicit heuristics, leverages stochastic simulations and adaptive tree expansion under uncertain environments, making it well-suited for large state spaces. It optimizes search results gradually through four steps: selection, expansion, simulation, and backpropagation, approaching the optimal solution. In contrast, A uses a heuristic function-guided deterministic search to guarantee optimal paths, but its performance depends on the design of the heuristic function. As a result, MCTS has been successfully applied to tasks such as RAG (Hu et al., 2024b; Jiang et al., 2024; Li and Ng, 2024; Feng et al., 2025), QA (Luo et al., 2025a; Gan et al., 2025), hallucinations mitigation (Cheng et al., 2025), text-to-SQL (Yuan et al., 2025b), etc. Additionally, Long (2023) trains an LLM controller using reinforcement learning to guide the LLM reasoner’s search path, and Chari et al. (2025) utilizes ant colony evolutionary algorithm to guide tree search.
@@
The value function evaluates the value of each state and guides the tree to expand towards branches with higher values in heuristic tree search. Xu (2023) train an energy function by noise-contrastive estimation as the value function. RAP (Hao et al., 2023) designs a series of heuristic value functions, including the likelihood of the action, the confidence of the state, self-evaluation results, and task-specific reward, and combines them according to task requirements. Reliable and generalized value functions facilitate the application of MCTS to more complex problems with deeper search spaces. AlphaMath (Chen et al., 2024b) and TS-LLM (Feng et al., 2024a) replace the hand-crafted value function with a learned LLM value function, automatically generating reasoning process and step-level evaluation signals in MCTS. VerifierQ (Qi et al., 2024b) integrates implicit Q-learning and contrastive Q-learning to train the value function, effectively mitigating the overestimation issue at the step level. Traditional MCTS methods expand only one trajectory, while rStar (Qi et al., 2024a) argues that the current value function struggles to guide the selection of the optimal path accurately. Therefore, rStar retains multiple candidate paths and performs reasoning with another LLM, ultimately selecting the path where both LLMs’ reasoning results are consistent. Gao et al. (2024c) propose SC-MCTS, which combines multiple reward models, including contrastive reward, likelihood, and self-evaluation as value functions. MCTSr (Zhang et al., 2024d) and SR-MCTS (Zhang et al., 2024g) take complete responses as nodes, expanding the search space through self-critique and correction. SR-MCTS utilizes pairwise preference rewards and global quantile score as the value function, offering a more robust value function estimation.
@@
Tree search can guide LLMs to generate long reasoning processes, and these data help train LLMs with stronger reasoning abilities (Zhai et al., 2024; Xu et al., 2024b; Guan et al., 2025). ReST-MCTS* (Zhang et al., 2024m) uses process rewards as a value function to guide MCTS, collecting high-quality reasoning trajectories and the value of each step to improve the policy model and reward model. Due to the step-by-step exploration of tree search, it can obtain finer-grained step-level feedback signals. MCTS-DPO (Xie et al., 2024) collects step-level preference data through MCTS and uses DPO for preference learning. AlphaLLM-CPL (Wang et al., 2024k) ranks trajectories based on preference reward gaps and policy prediction gaps, employing curriculum learning to efficiently utilize MCTS-collected trajectories. Recently, many LMRs (Qin et al., 2024a; Zhao et al., 2024b; Zhang et al., 2024h) have also confirmed the necessity of using tree search to construct high-quality long reasoning chain data for training.
@@
5. Advanced Topics
5.1. Generalizable System-2 Model
Currently, most LRMs exhibit strong reasoning performance in specific domains such as math and code, but they struggle to generalize to cross-domain, cross-lingual, or general tasks (Son et al., 2025; Zhao et al., 2025b). On the one hand, as the foundation of current System-2 models, CoT shows little effectiveness in non-symbolic reasoning tasks (Sprague et al., 2024). On the other hand, verifiers and critics have limited generalization capabilities, making it difficult to provide effective guidance to the reasoning model (LeVine et al., 2023; Kim et al., 2024a; Chen et al., 2024f). For the former limitation, integrating external tools and multi-agent feedback is a promising direction (Nathani et al., 2023; Lan et al., 2024). Many works aim to enable reasoning models to learn to perform retrieval or tool use during the reasoning process. Search-o1 (Li et al., 2025b) triggers external search tools by generating special tokens that represent search actions, and Search-R1 (Jin et al., 2025) further enhances the interaction between the LLM and search tools through RL. Deep Research (OpenAI, 2025; Google, 2024; Zheng et al., 2025; Li et al., 2025e) integrates various tools such as web search and code execution, demonstrating distinguished performance on general tasks like information synthesis, report writing, and complex code generation. For the latter challenge, some works utilize multi-objective training (Wang et al., 2024m), model ensemble (Lin et al., 2024c), soft reward (Su et al., 2025) or regularization constraints (Yang et al., 2024a; Jia, 2024) to make verifiers more generalizable. Compared to discriminative verifiers, generative critics inherit the generalization ability of LLMs and can leverage test-time compute to realize greater potential (Kim et al., 2025). GenPRM (Zhao et al., 2025a) and GRM (Liu et al., 2025e) leverage repeated sampling for test-time compute, and GRM trains a Meta RM to evaluate the quality of the critic’s outputs. RM-R1 (Chen et al., 2025b) and JudgeLRM (Chen et al., 2025a) use RL training to enable the critic with self-correction ability.